# Six-CD: Benchmarking Concept Removals for Benign Text-to-image Diffusion Models

### v1

@@ 5.2 Removal Ability on Multiple Concepts
In Sec. 5.1, we evaluate the ability to remove a single concept from one model. However, removing multiple concepts from one model is also important since the model builder usually needs to consider multiple malicious concepts in practice. In this subsection, we present the results of removing multiple concepts of celebrities and copyrighted characters in Fig. 6. In Fig. 6, the y-axis represents the detection rate, and a higher detection rate means a weaker concept removal performance. There are two observations from Fig. 6:
@@ 5.3 In-prompt and Out-prompt Retainability
In this subsection, we evaluate the generation ability on benign concepts using both in-prompt and out-prompt CLIP scores. In Fig. 7, we use DVD to test the in-prompt CLIP score and a subset with 500 benign prompts from LAION [41] for the out-prompt CLIP score. The observations as follows.
Observation 1: In-prompt retainability performs worse than out-prompt retainability. As shown in Fig. 7, the in-prompt CLIP score is lower than the out-prompt CLIP score across all the methods and concept categories. This indicates that the ability to generate benign parts in prompts containing malicious concepts is more negatively affected than the ability to generate totally benign prompts. As mentioned in Sec. 4, the in-prompt retainability is also important for concept removal methods. This inspires future research to pay more attention to the in-prompt retainability. Remark. It could be unfair to directly compare the two CLIP scores since they use different prompt sets to calculate CLIP scores. To further validate this observation, we use the clean version of DVD to calculate the out-prompt CLIP score, which is detailed in Appd. D.2. By comparing it with the in-prompt CLIP score in Appd. D.2, we have observed a consistent phenomenon with this subsection.
@@ 6 Conclusion and Ethical Statement
In this work, we address the lack of a systematic benchmark for concept removal methods by introducing a comprehensive and effective dataset and a new evaluation metric. Using this benchmark, we conduct an extensive evaluation of concept removal methods. Our experimental observations provide valuable and practical insights for future research in this field. While our benchmark may have a limitation in the inconvenient detection metric for art styles (using the CLIP score for art styles, instead of the detection rate from a classifier like other categories), it still offers a thorough and practical evaluation across the various settings of the methods.
Ethical statement. This work provides a standardized and comprehensive evaluation framework for concept removal methods, which are targeted to mitigate the potential malicious use of text-to-image diffusion models. Our research is conducted responsibly, transparently, and deeply dedicated to ethical standards. Despite involving the generation of sensitive content such as nudity and violence, it is strictly for research purposes and does not intend to produce or promote inappropriate material. On the contrary, our work aims to advance efforts to prevent the generation of inappropriate content.
@@ Appendix D Additional Experiments
D.1 Table of Removal Ability with Error Bar
Besides Fig. 4, we also report the results of removal ability and the error bar in Table 4. We calculate the standard variance using the estimation of bootstrap [44, 45].
@@ D.1 Table of Removal Ability with Error Bar
D.2 Out-prompt CLIP Score by Clean Version of DVD
@@ D.3 Time Cost
D.4 Fine-grained Retainability for Similar Concepts

### v2

@@ 5.2 Removal Ability on Multiple Concepts
In Sec. 5.1, we evaluate the ability to remove a single concept from one model. However, removing multiple concepts from one model is also important since the model builder usually needs to consider multiple malicious concepts in practice. In this subsection, we present the results of removing multiple concepts of celebrities and copyrighted characters in Fig. 6, where the y-axis represents the detection rate, and a higher detection rate means a weaker concept removal performance. We have two observations from Fig. 6.
@@ 5.3 In-prompt and Out-prompt Retainability
We evaluate the generation ability on benign concepts using both in-prompt and out-prompt CLIP scores. In Fig. 7, we use DVD to test the in-prompt CLIP score and a subset with benign prompts from LAION [37] to test out-prompt CLIP score. The observations as follows.
Observation 1: In-prompt retainability performs worse than out-prompt retainability. As shown in Fig. 7, the in-prompt CLIP score is lower than the out-prompt CLIP score across all the methods and concept categories. This indicates that the ability to generate benign parts in prompts containing malicious concepts is more negatively affected than the ability to generate totally benign prompts. As mentioned in Sec. 4, the in-prompt retainability is also important for concept removal methods. This inspires future research to pay more attention to the in-prompt retainability. Remark. It could be unfair to directly compare the two CLIP scores since they use different prompt sets to calculate CLIP scores. To further validate this observation, we use the clean version of DVD to calculate the out-prompt CLIP score, which is detailed in Appd. D.3. By comparing it with the in-prompt CLIP score in Appd. D.3, we have observed a consistent phenomenon with this subsection.
@@ 6 Conclusion and Ethical Statement
In this work, we address the lack of a systematic benchmark for concept removal methods by introducing a comprehensive and effective dataset and a new evaluation metric. Using this benchmark, we conduct an extensive evaluation of concept removal methods. Our experimental observations provide valuable and practical insights for future research in this field. While our benchmark may have a limitation in the inconvenient detection metric for art styles (using the CLIP score for art styles, instead of the detection rate from a classifier like other categories), it still offers a thorough and practical evaluation across the various method settings.
Ethical statement. This work provides a comprehensive evaluation framework for concept removal, which is targeted to mitigate the potential malicious use of text-to-image diffusion models. Our research is conducted responsibly, transparently, and deeply dedicated to ethical standards. Despite involving the generation of sensitive content such as nudity and violence, it is strictly for research purposes and does not intend to produce or promote inappropriate material. On the contrary, our work aims to advance efforts to prevent the generation of inappropriate content.
Acknowledgement
Jie Ren, Kangrui Chen, Yingqian Cui, Shenglai Zeng, and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IOS2107215, DUE2234015, CNS2246050,
DRL2405483 and IOS2035472, the Army Research Office (ARO) under grant number W911NF-21-1-0198, Amazon Faculty Award, JP Morgan Faculty Award, Meta, Microsoft and SNAP.
Supplementary Material
@@ Appendix D Additional Experiments
D.1 Human Evaluation of Proposed New Metric
We provide additional human evaluations Table 5 to demonstrate the validity of the proposed metric (in-prompt CLIP score).
Settings: We evaluate the score based on 50 pairs of images. Each pair of images are generated using the same malicious prompt and using two different concept removal methods. In addition to the malicious prompt, we use the corresponding clean version of the malicious prompt in Dual-Version Dataset. Given the pair of images and the clean prompt, we ask humans to rank which image better aligns with the clean prompt. After collecting the ranking results from 15 humans, we compare the ranking with the order given by the in-prompt CLIP scores of the two images. If the human preference is the same as the CLIP score, we label the sample as correct. With 15 humans, there are a total of 750 comparisons done in our evaluation.
Results: We report the accuracy in the following table. Note that if the two CLIP scores are close to each other, the human preference may be inaccurate. Therefore, we report the results when the difference is larger than a threshold $T$. From the table below, we can see that when $T=0$, the accuracy is 86.86%, which means most of the in-prompt CLIP score is consistent with human preference. When $T=0.03$, the correctness is even larger than 95%. These observations validate the effectiveness of our proposed metric.
Table 5: Human Evaluation of in-prompt CLIP score
$T$ 0 0.025 0.03
Acc. 86.86% 91.24% 95.24%
D.2 Table of Removal Ability with Error Bar
Besides Fig. 4, we also report the results of removal ability and the error bar in Table 6. We calculate the standard variance using the estimation of bootstrap [^2] https://wires.onlinelibrary.wiley.com/doi/full/10.1002/wics.182, https://ieeexplore.ieee.org/abstract/document/4767957, https://yuleii.github.io/2021/01/22/bootstrap.html.
@@ D.2 Table of Removal Ability with Error Bar
D.3 Out-prompt CLIP Score by Clean Version of DVD
@@ D.4 Time Cost
D.5 Fine-grained Retainability for Similar Concepts
@@ D.6 FID
D.7 Experiments on other models
Besides SD v1.4, we provide additional experiments on DPO-Diffusion and SD v1.5 in this section. The observations in this section are consistent with the findings in our paper:
  • General concepts are harder to remove. In the following Table 8, we choose Nudity and Copyright to represent general and specific concepts, respectively. Similar to our findings in Section 5.1, after concept removals, the generated general concepts are higher than specific concepts in different methods.
  • Inference-time methods fail in removing multiple concepts. In the following Table 9, the two inference-time methods, NEG and SLD, have poor ability in removing unwanted concepts when the number of concepts increases to 100. To explain this, as mentioned in Section 5.2, these methods have to encode the string containing all the concepts in the embeddings of one single prompt. The long string will exceed the capacity of the text encoder of T2I diffusion models and lead to failed removal.
  • Closed-form solutions by modifying linear components perform well in removing multiple concepts. As shown in the following Table 9, UCE performs well in removing 100 concepts: only 0.75% of images are detected with unwanted concepts, while NEG is 81.05%. We conjecture that combining and changing multiple concepts in the linear components is easier than in other non-linear parts.
  • In-prompt retainability performs worse than out-prompt retainability. In the following Table 10, the in-prompt CLIP score is lower than the out-prompt CLIP score across different methods and concept categories. This result is also consistent with our observation in Section 5.3.
Table 8: Detection rate of nudity and copyright concepts after concept removals (model: DPO-Diffusion).
no CR NEG EMCID SDD UCE
Nudity 0.8960 0.3258 0.8937 0.0828 0.1482
Copy. 0.9736 0.2560 0.3088 0.0048 0.0152
Table 9: Detection rate after removing multiple concepts (model: SD v1.5).
Copyright number NEG SLD UCE
1 0.2090 0.0816 0.0144
100 0.8105 0.7530 0.0075
Table 10: CLIP score (model: DPO-Diffusion)
EMCID SDD
Concept category nudity copyright object nudity copyright object
Out-prompt retainability 0.3320 0.3313 0.3314 0.3190 0.3129 0.2601
In-prompt retainability 0.2929 0.2400 0.2838 0.2625 0.2388 0.2277
Figure 12: Effectiveness on SDXL and SD3.
Table 11: Detection Rate and Retainability on SD3
Detect. Harm Nudity Celeb. Copy. Obj. Art
ORG 0.637 0.174 0.980 1.000 0.980 0.275
NEG 0.467 0.071 0.143 0.714 0.786 0.244
ESD 0.628 0.138 0.393 1.000 0.643 0.246
SPM 0.633 0.180 0.143 1.000 0.393 0.268
SDD 0.635 0.131 0.679 0.321 0.214 0.273
SLD 0.495 0.107 0.143 0.964 0.929 0.255
Retain. Harm Nudity Celeb. Copy. Obj. Art
NEG 0.274 0.286 0.273 0.271 0.298 0.314
ESD 0.274 0.290 0.264 0.259 0.277 0.315
SPM 0.260 0.290 0.248 0.252 0.269 0.315
SDD 0.258 0.291 0.265 0.250 0.240 0.312
SLD 0.273 0.290 0.266 0.264 0.296 0.315
We also discuss the transferability of Six-CD and DVD as follows.
(1) Prompt effectiveness of Six-CD. In Fig. 12, we show the effectiveness (i.e. $n/N$) of Harm and Nudity on SDXL and SD3. While our dataset generates fewer unwanted concepts on these models, it remains much more effective than other datasets like I2P. Specifically, its effectiveness is around three times as high as I2P for Nudity and twice for Harm.
(2) Results of removal and retainability. In Table 11, we provide results of 5 removal methods on SD3. We observe similar trends as in SD1. For example, prompts of specific concepts have higher effectiveness than general concepts. The detection of specific unwanted concepts are higher than 0.98 on ORG (i.e. models without concept removal). Another key observation, consistent with SD1, is that general concepts are more difficult to remove. Specifically, on general concepts, concept removal methods result in only a minimal reduction in detection rates compared to ORG.